# Cyclic Vision-Language Manipulator: Towards Reliable and Fine-Grained Image Interpretation for Automated Report Generation

### v2

@@ Abstract
Despite significant advancements in automated report generation, the opaqueness of text interpretability continues to cast doubt on the reliability of the content produced. This paper introduces a novel approach to identify specific image features in X-ray images that influence the outputs of report generation models. Specifically, we propose Cyclic Vision-Language Manipulator CVLM, a module to generate a manipulated X-ray from an original X-ray and its report from a designated report generator. The essence of CVLM is that cycling manipulated X-rays to the report generator produces altered reports aligned with the alterations pre-injected into the reports for X-ray generation, achieving the term “cyclic manipulation”. This process allows direct comparison between original and manipulated X-rays, clarifying the critical image features driving changes in reports and enabling model users to assess the reliability of the generated texts. Empirical evaluations demonstrate that CVLM can identify more precise and reliable features compared to existing explanation methods, significantly enhancing the transparency and applicability of AI-generated reports.
@@ 1 Introduction
In response, several studies have adopted existing Explainable AI (XAI) techniques to uncover the visual features influencing generated content, thereby aiming to enhance the trustworthiness of the generated texts. The most widely used XAI methods in this field typically generate an attribution map to highlight the influential features related to the generated contents through cross-attention maps Liu et al. (2019); Cao et al. (2023); Chen et al. (2020) or the GradCAM-based methods Alfarghaly et al. (2021); Spinks and Moens (2019); Wang et al. (2024). However, these attribution maps often fail to precisely highlight fine-grained visual features, instead showing large and coarse areas. More critically, there is currently no reliable method to verify whether the attribution maps accurately reflect the information utilized by report generators to produce findings, as there is no access to the ground truth regarding what the model has actually learned Chase Walker et al. (2024). However, attribution maps often fail to highlight fine-grained visual features, instead presenting large and coarse areas (see Fig. 1 (A)). More critically, there is currently no reliable method to verify whether these attribution maps accurately represent the information utilized by report generators in producing findings, as there is no ground truth to confirm what the model has actually learned Chase Walker et al. (2024). These two limitations raise an important ambiguity: while the highlighted areas in the attribution maps may appear irrelevant to the generated findings, it remains unclear whether report generators rely on irrelevant features or whether the explanation methods misidentify visual features relative to the content being explained.
@@
To address these limitations and provide a reliable attribution map that can facilitate the understanding of generated texts, we propose a model-agnostic counterfactual explanation method to generate attribution maps. This method relies on the creation of counterfactual images, which modify specific features in the images to elicit changes in the model’s decision. Comparing the counterfactual image to the original enables the identification of crucial features responsible for the altered decision Wachter et al. (2017). Specifically, we introduce a module called the Cyclic Vision-Language Manipulator (CVLM) to generate effective counterfactual images using a text-conditioned diffusion model. The core of CVLM is to achieve cyclic manipulation: the manipulation begins with the generated text, and the counterfactual images derived from this manipulated text must produce a regenerated report that accurately reflects the same alterations introduced in the initial report (see Fig. 2 (B)). The comparison between the cyclically generated counterfactual image and the original X-ray produces a difference map to identify fine-grained image features. Furthermore, the counterfactual images, coupled with the altered regenerated report, verify that the identified features are being utilized by existing report generators in their text generation process. The contributions of our work can be summarized as follows:
@@
  • Experiments demonstrate that the model-agnostic nature of CVLM allows users to evaluate whether the visual features associated with the generated findings are meaningful for the corresponding content. This capability also enables users to compare the reliability of results across different report generators and make informed decisions on adopting suggestions from various models.
@@ 2.1 Counterfactual explanation
The most widely applied methods for visual explanation are post-hoc explanation methods, as their model-agnostic nature enables them to generalize across different models. Popular approaches include activation-based methods, backpropagation-based methods, and perturbation-based methods. With the recent evolution of generative AI models, counterfactual explanation methods have emerged as an advanced perturbation-based approach. These methods generate counterfactual images that elicit different findings from the model, enabling users to identify differences between similar classes—a common challenge in medical image classification tasks. Applications has emerged in explaining the classifiers with different modalities in medical image analysis, including X-rays Atad et al. (2022); Mertes et al. (2022); Singla et al. (2023); Schutte et al. (2021); Sankaranarayanan et al. (2022), Magnetic Resonance Imaging Tanyel et al. (2023); Fontanella et al. (2023), ultrasound Reynaud et al. (2022), histopathology images Karras et al. (2020); Schutte et al. (2021), and Computed Tomography Fang et al. (2024a, b). Over time, counterfactual generation methods have evolved from using variational autoencoders Rodriguez et al. (2021) and generative adversarial networks Lang et al. (2021); Atad et al. (2022) to leveraging diffusion models Rombach et al. (2022); Fang et al. (2024a, b). While these methods typically generate counterfactual images to elicit changes under the guidance of a given classifier, in this work, we aim to develop a counterfactual generation method driven by text manipulation. The counterfactual X-ray is designed to achieve the desired content changes in the regenerated reports.
@@ 2.2 Explainability in report generation models
In contrast to these approaches, Tanida et al. (2023) proposed a report generation method called RGRG, which significantly enhances the interpretability and transparency of generated reports by tailoring content generation to specific anatomical regions. However, this approach heavily relies on the preparation of a large paired dataset comprising anatomical regions and corresponding fine-grained reports for both the anatomical detection model and the report generation training. This requirement substantially increases manual labeling costs and limits the model’s ability to incorporate larger X-ray-report datasets that lack fine-grained annotations. Additionally, the self-explanatory workflow of RGRG is not transferable to other report generators with different generation processes. In this paper, we aim to propose an explanation method that achieves fine-grained visual feature localization for generated key findings as RGRG, while eliminating the need for extensive labeling. Being model-agnostic, our approach can enhance the interpretability and transparency of generated reports across a wide range of report generation models.
@@ 3 Method
The overall framework for CVLM and applying it for decoding a target report generator is illustrated in Fig. 3. In the following, we detail the components and application of CVLM to obtain tailored report explanations for a target report generator.
Figure 3: Overview of applying the proposed Cyclic Vision-Language Manipulator (CVLM) in explaining the report generator and the challenges existing in developing the CVLM.
@@ 3.1 Development of CVLM
3.1.1 Data prepration
To enable image manipulation based on the image features learned for the generated keywords by the report generator, rather than relying on real image features corresponding to medical terminology, we inferred the generated reports from the target report generator on the dataset it was trained on. We paired the initial X-ray image with the inferred results and trained the model to reconstruct the initial X-ray using the generated reports instead of the human-labeled ground-truth (GT) reports. This model is referred to as the tailored generation model. The tailored model can generate counterfactual X-rays by removing keywords from the generated reports, enabling the detection of specific differences responsible for changes in the findings produced by the target report generator.
@@ I R M
where $z_{T}$ is the encoded feature of the initial query X-ray image from the encoder of a variational autoencoder, $x_{0}$, added with a Gaussian noise $\epsilon$, $\tau_{\theta}$ is the text encoder than transforms the prompt to the text embedding. During our training, we leverage the pretrained model weight for the text embedding and image autoencoder modules by a stable diffusion model pretrained on MIMIC Liang et al. (2023). During training, we initialize the weight of UNET architecture by the stable diffusion pretrained weight “CompVis/stable-diffusion-v1-4” and freeze the parameter in the image autoencoder.
@@ 3.2.2 Unsupervised feature identification frame generation
Figure 5: Pipeline of the unsupervised frame generation method.
@@ 4.1.2 Implementation details
For developing the CVLM, we frozen the text encoders and trained the Diffusion using a batch size of 8 and a learning rate of 5e-5 on one A6000 GPU with 40 GB of memory. We trained the model with 100k steps over about one week. The final model for cyclic counterfactual generation was selected based on the highest PSNR of reconstructed image achieved on the validation set. For counterfactual generation, the DDIM step is set to 50. For the frame mask generation, the Gaussian blur is set at $5\times 5$, the threshold is set between $95\pm 10$ which we select the best value selected for each manipulated finding, keeping $K$ at 5. More details are given in Appendix 1. Most of the experiments were conducted on the A6000 GPU, while those reported in Appendix Table 2.3 were run on an RTX 8000 GPU with 48 GB of memory.
@@ 4.1.3 Evaluation methods
We conduct a thorough evaluation of the proposed method in the following four steps: (1) We first propose a metric called Cyclic Rate (CR) to quantitatively assess the effectiveness of CVLM in generating cyclic counterfactual images for explanation. CR is defined as the success rate of achieving the intended manipulation in the regenerated report. Specifically, CR is calculated as the ratio of counterfactual images generated from reports where a finding is removed, and whose regenerated report successfully reflects the removal of that finding, to the total number of manipulations performed; (2) Next, we apply the feature identification framework to decode the visual features utilized by two individual report generators and compare the features they utilize; (3) To further illustrate the superiority of CVLM in identifying fine-grained and verifiable features for generated contents, we compare our proposed difference detection frame (introduced in Section 3.2) with one most popular model-agnosic explanation method Cross-Attention Map and the explanation results from a state-of-the-art self-explainable report generator RGRG Tanida et al. (2023). The results are shown in Fig. 8. (4) Finally, we conduct ablation studies on the Cyclic Rate of CVLM, analyzing its performance with respect to inference times ($T$), training durations, and the use of raw reports versus structured reports.
@@ 4.2.1 Success rate of cyclic counterfactual generation
Table. 1 shows the quantitative results of CVLM in obtaining the cyclic counterfactual examples for R2Gen and R2GenCMN respectively. The success rate is calculated by running a script to manipulate 569 findings generated from 400 images and evaluating the counterfactual images based on their ability to alter the regenerated reports in accordance with the applied manipulations. Both models achieve a success rate around 0.7 with CVLM for R2GEN achieving a higher manipulation success rate.
4.2.2 Comparison of different report generators
As “Atelectasis” is the point of disagreement between the two models, we included the opinions of two radiologists to provide their judgment and identify the features indicative of its presence. We found that their assessments and the identified features closely matched the features detected by R2GenCMN. This suggests that R2GenCMN utilizes clinically recognized features for generating this finding, even though the ground-truth report from the dataset did not include it. This also indicates that the explanations provided for the generated reports not only validate the findings of the models but also help identify potential missing findings in reports produced by humans. By presenting the generated report alongside its feature maps, our approach offers a more reliable and safer method for utilizing the generated reports in clinical practice. More cases are given in Appendix 2.1.
Figure 6: Report findings of R2GenCMN: Cardiomegaly, Support Devices, Atelectasis
Figure 7: Report findings of R2Gen: Enlarged Cardiomediastinum, Support Devices
Support Device
Refer to captionRefer to captionRefer to captionRefer to caption
‘endotracheal tube’
Cardiomegaly
Refer to captionRefer to captionRefer to captionRefer to caption
‘heart size’
Enlarged Cardiomediastinum
Refer to captionRefer to captionRefer to captionRefer to caption
‘cardiomediastinal silhouette’
Lung Opacity
Refer to captionRefer to captionRefer to captionRefer to caption
‘opacities’
Edema
Refer to captionRefer to captionRefer to captionRefer to caption
‘vascular congestion’
Consolidation
Refer to captionRefer to captionRefer to captionRefer to caption
‘consolidation’
Figure 9: Qualitative comparison with cross attention and RGRG methods on the MIMIC-CXR dataset. 1st row is the initial images, 2nd row is the counterfactuals generated by CVLM method, 3rd row is the images with the bounding box and the text generated by RGRG method, and 4th row is the heatmap with the attention entities (the blue text) generated by the cross attention method. Note: The counterfactuals in the figure all achieve the cyclic manipulation in the regenerated report. In this image, CVLM and attention are model-agnostic explanation methods applied to the findings in the generated report from R2GenCMN, which are confirmed to exist in the GT report. In contrast, RGRG generates its report separately based on specific regions. Please zoom in for detailed observation.
4.2.3 Benchmark against state-of-the-art explanation methods
Additionally, CVLM not only demonstrates the ability to identify fine-grained features within X-rays but also effectively manipulates features occurring in multiple locations, such as opacity. It can also detect large-scale global abnormalities, such as enlarged cardiomediastinum, significantly surpassing the feature identification capabilities of existing explanation methods. Additional examples are provided in Appendix 2.2.
4.2.4 Ablation study
Table 1: Ablation study on the training iterations of the image generation model for CVLM in explaining the generator models R2GenCMN and R2Gen.
@@ 4.2.4 Ablation study
We investigated the impact of training time on the stable diffusion model for achieving CVLM. Specifically, we compared models achieving the best reconstruction (highest PSNR) and models trained with additional iterations. Table 1 demonstrates that the model with the best reconstruction ability, when paired with the generated text, achieves the highest cyclic manipulation effectiveness for explaining the report.
Additionally, we compared the manipulation method within CVLM to a direct report manipulation approach, where Stable Diffusion is trained directly with unprocessed reports without pre-cleaning. The results in Fig. 10 indicate that using organized prompts focused on findings results in significantly better performance compared to removing entire sentences from unstructured reports. For quantitative comparison, out of a total of 400 manipulations, the structured report manipulation achieved a much higher success rate (0.712) than the direct report manipulation approach (0.449) for the R2Gen report generator.
Finally, we analyzed the influence of the denoising step $T$ during diffusion model inference on achieving cyclic success. We evaluated $T$ values ranging from 20 to 100 and found that $T>50$ yields the highest success rates. For efficiency, $T=50$ was selected. The computation cost and numerical results are provided in Appendix 2.3.
(a) Remove “cardiomegaly” from the raw report.
@@
5 Discussion
5.1 Clinical significance.
The major contribution of this paper is the development of a method capable of identifying feature attributions that can be verified as being used by the report generator for its generated content. The identified features may be consistent or inconsistent with human knowledge, as the absence of “ground truth” for the features learned by the model. The verified features identified by our method can assist human experts who use report generators in several ways: they can help users locate relevant features for hard-to-detect abnormalities and evaluate the reliability of the findings from the automated report generator by comparing the identified features with their own knowledge.
5.2 Broader applications.

### v3

@@ Abstract
Despite significant advancements in automated report generation, the opaqueness of text interpretability continues to cast doubt on the reliability of the content produced. This paper introduces a novel approach to identify specific image features in X-ray images that influence the outputs of report generation models. Specifically, we propose Cyclic Vision-Language Manipulator (CVLM), a module to generate a manipulated X-ray from an original X-ray and its report from a designated report generator. The essence of CVLM is that cycling manipulated X-rays to the report generator produces altered reports aligned with the alterations pre-injected into the reports for X-ray generation, achieving the term “cyclic manipulation”. This process allows direct comparison between original and manipulated X-rays, clarifying the critical image features driving changes in reports and enabling model users to assess the reliability of the generated texts. Empirical evaluations demonstrate that CVLM can identify more precise and reliable features compared to existing explanation methods, significantly enhancing the transparency and applicability of AI-generated reports.
$1$The appendix for this work is available at https://arxiv.org/abs/2411.05261.
@@ 1 Introduction
In response, several studies have adopted existing Explainable AI (XAI) techniques to uncover the visual features influencing generated content, thereby aiming to enhance the trustworthiness of the generated texts. The most widely used XAI methods in this field typically generate an attribution map to highlight the influential features related to the generated contents through cross-attention maps Liu et al. (2019); Cao et al. (2023); Chen et al. (2020) or the GradCAM-based methods Alfarghaly et al. (2021); Spinks and Moens (2019); Wang et al. (2024). However, these attribution maps often fail to highlight fine-grained visual features, instead presenting large and coarse areas (see Fig. 1 (A)). More critically, there is currently no reliable method to verify whether these attribution maps accurately represent the information utilized by report generators in producing findings, as there is no ground truth to confirm what the model has actually learned Chase Walker et al. (2024). These two limitations raise an important ambiguity: while the highlighted areas in the attribution maps may appear irrelevant to the generated findings, it remains unclear whether report generators rely on irrelevant features or whether the explanation methods misidentify visual features relative to the content being explained.
@@
To address these limitations and provide a reliable attribution map that can facilitate the understanding of generated texts, we propose a model-agnostic counterfactual explanation method to generate attribution maps. This method relies on the creation of counterfactual images, which modify specific features in the images to elicit changes in the model’s decision. Comparing the counterfactual image to the original enables the identification of crucial features responsible for the altered decision Wachter et al. (2017). Specifically, we introduce a module called the Cyclic Vision-Language Manipulator (CVLM) to generate effective counterfactual images using a text-conditioned diffusion model. The core of CVLM is to achieve cyclic manipulation: the manipulation begins with the generated text, and the counterfactual images derived from this manipulated text must produce a regenerated report that accurately reflects the same alterations introduced in the initial report (see Fig. 2 (B)). The comparison between the cyclically generated counterfactual image and the original X-ray produces a difference map to identify fine-grained image features. Furthermore, the counterfactual images, coupled with the altered regenerated report, verify that the identified features are being utilized by existing report generators in their text generation process. The major contribution of this paper is the development of a method capable of identifying feature attributions that can be verified as being used by the report generator for its generated content. The identified features may be consistent or inconsistent with human knowledge, as the absence of “ground truth” for the features learned by the model. The verified features identified by our method can assist model developers in locating relevant visual cues used by the generator, and help human experts evaluate the reliability of the generated findings by comparing these features with their own clinical knowledge. The key contributions of this work are summarized as follows:
@@
  • Experiments indicate that the model-agnostic nature of CVLM allows users to evaluate whether the visual features associated with the generated findings are meaningful for the corresponding content. This capability also enables users to compare the reliability of results across different report generators and make informed decisions on adopting suggestions from various models.
@@ 2.1 Counterfactual Explanation
The most widely applied methods for visual explanation are post-hoc explanation methods, as their model-agnostic nature enables them to generalize across different models. Popular approaches include activation-based methods, backpropagation-based methods, and perturbation-based methods. With the recent evolution of generative AI models, counterfactual explanation methods have emerged as an advanced perturbation-based approach. These methods generate counterfactual images that elicit different findings from the model, enabling users to identify differences between similar classes—a common challenge in medical image classification tasks. Applications has emerged in explaining the classifiers with different modalities in medical image analysis, including X-rays Atad et al. (2022); Mertes et al. (2022); Singla et al. (2023); Schutte et al. (2021); Sankaranarayanan et al. (2022); Maksudov et al. (2025), Magnetic Resonance Imaging Tanyel et al. (2023); Fontanella et al. (2023), ultrasound Reynaud et al. (2022), histopathology images Karras et al. (2020); Schutte et al. (2021), and Computed Tomography Fang et al. (2024a, b). Over time, counterfactual generation methods have evolved from using variational autoencoders Rodriguez et al. (2021) and generative adversarial networks Lang et al. (2021); Atad et al. (2022) to leveraging diffusion models Rombach et al. (2022); Fang et al. (2024a, b). While these methods typically generate counterfactual images to elicit changes under the guidance of a given classifier, in this work, we aim to develop a counterfactual generation method driven by text manipulation. The counterfactual X-ray is designed to achieve the desired content changes in the regenerated reports.
@@ 2.2 Explainability in Report Generation Models
Figure 3: Overview of applying the proposed Cyclic Vision-Language Manipulator in explaining the report generator and the challenges existing in developing the CVLM.
@@
In contrast to these approaches, Tanida et al. (2023) proposed a report generation method called RGRG, which significantly enhances the interpretability and transparency of generated reports by tailoring content generation to specific anatomical regions. However, this approach heavily relies on the preparation of a large paired dataset comprising anatomical regions and corresponding fine-grained reports for both the anatomical detection model and the report generation training. This requirement substantially increases manual labeling costs and limits the model’s ability to incorporate larger X-ray-report datasets that lack fine-grained annotations. Additionally, the self-explanatory workflow of RGRG is not transferable to other report generators with different generation processes.
In this paper, we aim to propose an explanation method that achieves fine-grained visual feature localization for generated key findings as RGRG, while eliminating the need for extensive labeling. Being model-agnostic, our approach can enhance the interpretability and transparency of generated reports across a wide range of report generation models.
@@ 3 Method
The overall framework for CVLM and applying it for decoding a target report generator is illustrated in Fig. 3. In the following, we detail the components and application of CVLM to obtain tailored explanations for a target report generator.
@@ 3.1 Development of CVLM
3.1.1 Data Preparation
To enable image manipulation based on the image features learned for the keywords generated by the report generator, rather than relying on real image features corresponding to medical terminology, we inferred the generated reports from the target report generator on the dataset on which it was trained. We paired the initial X-rays with the inferred results and trained the model to reconstruct the initial X-rays using the generated reports instead of the human-labeled ground-truth (GT) reports. This model is referred to as the tailored generation model, which can generate counterfactual X-rays by removing keywords from the generated reports, enabling the detection of specific differences responsible for changes in the findings produced by the target report generator.
@@ 3.1.2 Training Objective
where $z_{T}$ is the encoded feature of the initial real query X-ray image from the encoder of a variational autoencoder, $\tau_{\theta}$ is the text encoder that transforms the prompt to the text embedding. During our training, we leverage the pretrained model weight for the text embedding and image autoencoder modules by a stable diffusion model pretrained on MIMIC Liang et al. (2023). During training, we initialize the weight of UNET architecture by the stable diffusion pretrained weight “CompVis/stable-diffusion-v1-4” and freeze the parameter in the image autoencoder.
@@ 3.2.2 Unsupervised Feature Identification Frame Generation
Original
Manipulated
Diff (Filtered)
Opening
Dilation
Components
Top Contour
Frame
Figure 5: Image processing pipeline in chest X-ray analysis.
Query
Recon
Cardiomegaly
Device
Atelectasis
(a) R2GenCMN: Cardiomegaly, Support Devices, Atelectasis
(b) R2Gen: Enlarged Cardiomediastinum, Support Devices
@@ 4.1.2 Implementation Details
For developing CVLM, we froze the text encoders and trained the diffusion model using a batch size of 8 and a learning rate of $5\times 10^{-5}$ on a single A6000 GPU with 40 GB of memory. The model was trained for 100k steps over approximately one week. The final model for cyclic counterfactual generation was selected based on the highest PSNR of the reconstructed images on the validation set. For counterfactual generation, the DDIM step was set to 50. For frame mask generation, a Gaussian blur of size $5\times 5$ was applied, and the threshold was set to $95\pm 10$, with the best value selected for each manipulated finding. The number of preserved tokens, $K$, was fixed at 5. More details are provided in the online appendix [^1] The appendix for this work is available at https://arxiv.org/abs/2411.05261..
@@ 4.1.3 Evaluation Methods
We conduct a thorough evaluation of the proposed method in the following four steps: (1) We first propose a metric called Cyclic Rate (CR) to quantitatively assess the effectiveness of CVLM in generating cyclic counterfactual images for explanation. CR is defined as the success rate of achieving the intended manipulation in the regenerated report. Specifically, CR is calculated as the ratio of counterfactual images generated from reports where a finding is removed, and whose regenerated report successfully reflects the removal of that finding, to the total number of manipulations performed; (2) Next, we apply the feature identification framework to decode the visual features utilized by two individual report generators and compare the features they utilize; (3) To further illustrate the superiority of CVLM in identifying fine-grained and verifiable features for generated contents, we compare our proposed difference detection frame (introduced in Section 3.2) with a widely used model-agnosic explanation method Cross-Attention Map and the explanation results from a state-of-the-art self-explainable report generator RGRG Tanida et al. (2023). The results are shown in Fig. 6. (4) Finally, we conduct ablation studies on the Cyclic Rate of CVLM, analyzing its performance with respect to inference times ($T$), training durations, and the use of raw reports versus structured reports.
@@ 4.2.1 Success Rate of Cyclic Counterfactual Generation
In Table 1, we first investigate the impact of different training durations in the cyclic rate (CR) schedule of CVLM on generating cyclic counterfactual explanations. We then compare the CR settings for R2Gen and R2GenCMN. Specifically, we evaluate models that achieve the highest reconstruction quality (measured by PSNR) and those that achieve the lowest FID score. Our results show that the model with the best reconstruction ability, when paired with the generated text, leads to the highest cyclic manipulation effectiveness in explaining the reports for both R2Gen and R2GenCMN. We therefore select this checkpoint for report explanation. Both models achieve a success rate around 0.7, with CVLM for R2Gen showing a higher manipulation success rate overall.
@@
Table 1: Success rate of CVLM in explaining R2GenCMN and R2Gen across training checkpoints selected by PSNR and FID. Success rates are calculated over 569 findings from 400 images. 14k and 16k denote best-PSNR checkpoints; other iterations are selected based on lowest FID.
Support DeviceRefer to captionRefer to captionRefer to captionRefer to caption ‘endotracheal tube’ CardiomegalyRefer to captionRefer to captionRefer to captionRefer to caption ‘heart size’ CardiomediastinumRefer to captionRefer to captionRefer to captionRefer to caption ‘cardiacsilhouette’ EdemaRefer to captionRefer to captionRefer to captionRefer to caption ‘vascular congestion’ Lung OpacityRefer to captionRefer to captionRefer to captionRefer to caption ‘opacities’ ConsolidationRefer to captionRefer to captionRefer to captionRefer to caption ‘consolidation’
Figure 7: Qualitative comparison with cross attention and RGRG methods on the MIMIC-CXR dataset. 1st row is the initial images, 2nd row is the counterfactuals generated by CVLM method, 3rd row is the images with the bounding box and the text generated by RGRG method, and 4th row is the heatmap with the attention entities (the blue text) generated by the cross attention method. Note: The counterfactuals in the figure all achieve the cyclic manipulation in the regenerated report. In this image, CVLM and attention are model-agnostic explanation methods applied to the findings in the generated reports from R2GenCMN, which are confirmed to exist in the GT report. In contrast, RGRG are explainable report generator which generates its report separately based on specific regions.
(a) Remove the “cardiomegaly” sentence from the raw report.
@@
4.2.2 Comparison of Different Report Generators
As “Atelectasis” is the point of disagreement between the two models, we included the opinions of two radiologists to provide their judgment and identify the features indicative of its presence. We found that their assessments and the identified features closely matched the features detected by R2GenCMN. This suggests that R2GenCMN utilizes clinically recognized features for generating this finding, even though the ground-truth report did not include it. This also indicates that the explanations provided for the generated reports not only validate the findings of the models but also help identify potential missing findings in reports produced by humans. By presenting the generated report alongside its feature maps, our approach offers a more reliable and safer method for utilizing the generated reports in clinical practice. More cases are given in Appendix B [^1] The appendix for this work is available at https://arxiv.org/abs/2411.05261..
4.2.3 Benchmark against State-of-the-art Explanations
Additionally, CVLM not only demonstrates the ability to identify fine-grained features within X-rays but also effectively manipulates features occurring in multiple locations, such as opacity. It can also detect large-scale global abnormalities, such as enlarged cardiomediastinum, significantly surpassing the feature identification capabilities of existing explanation methods.
4.2.4 Ablation Study
We compared the manipulation method within CVLM to a direct report manipulation approach, where Stable Diffusion is trained directly with unprocessed reports without pre-cleaning. The results in Fig. 8 indicate that using organized prompts focused on findings results in significantly better performance compared to removing entire sentences from unstructured reports. For quantitative comparison, out of a total of 400 manipulations, the structured report manipulation achieved a much higher success rate (0.712) than the direct report manipulation approach (0.449) for the R2Gen model.
Finally, we analyzed the influence of the denoising step $T$ during diffusion model inference on achieving cyclic success. We evaluated $T$ values ranging from 20 to 100 and found that $T>50$ yields the highest success rates. For efficiency, $T=50$ was selected. The computation cost and numerical results are provided in Appendix B [^1] The appendix for this work is available at https://arxiv.org/abs/2411.05261..
4.3 Discussion
4.3.1 Broader Applications